# Continual Learning of Large Language Models: A Comprehensive Survey

### v2

@@ 2.2.2. Techniques of Continual Learning
The objectives for past domains are often challenging to measure or optimize due to the memory constraints (Definition 2.3). Therefore, the core of designing CL algorithms lies in identifying a proxy learning objective for the first term without violating the memory constraint. Existing CL techniques can be roughly categorized into 5 groups: (i) replay-based, (ii) regularization-based, (iii) architecture-based, (iv) optimization-based, and (v) representation-based (de2021continual, ; wang2024comprehensive, ). Here, we provide a concise yet comprehensive introduction to the first three categories of continual learning techniques, as they find extensive application in continually learning large language models.
where ^$\widehat{{\mathcal{L}}}_{S}$ denotes the empirical loss term evaluated on the set of examples $S$. Often regarded as a simplistic solution to CL, replay-based methods may theoretically lead to loose generalization bounds (shi2024unified, ). Despite this, they are valued for their simplicity, stability, and high performance, even with a small episodic memory (chaudhry2019tiny, ; riemer2018learning, ). For instance, DER++ (buzzega2020dark, ) demonstrates consistent performance enhancement by replaying a small set of past examples along with their logits (known as dark experience replay). ESM-ER (sarfraz2023error, ) introduces error sensitivity modulation (ESM) to mitigate abrupt representational drift caused by high-error new examples. A significant focus in replay-based CL is enhancing sample efficiency for buffer maintenance. For instance,  (rebuffi2017icarl, ) prioritizes exemplar selection based on herding to accurately model class mean throughout class-incremental learning. (zhao2022memory, ) propose storing low-fidelity examples to achieve memory-efficient exemplar set maintenance. RM (Rainbow Memory) (bang2021rainbow, ) introduces diversity-aware memory updates based on per-sample uncertainty estimation and data augmentation for class-incremental learning.
@@
Architecture-Based Methods. Expanding the network architecture dynamically to assimilate new knowledge is deemed the most efficient form of CL (wang2022learning, ; wang2022dualprompt, ). This method primarily tackles adaptation challenges and can achieve zero-forgetting when task IDs are available during inference or can be correctly inferred (gururangan2022demix, ; wistuba2023, ). However, due to the difficulty of task ID inference, architecture expansion is predominantly utilized in TIL but is scarcely explored in DIL or CIL. Progressive Neural Networks (PNN) (rusu2016progressive, ) proposes learning laterally connected neurons as new tasks arise, ensuring non-forgetting and enabling transfer of previously learned neurons for future tasks. In conjunction with pre-trained backbone large models like ViT (dosovitskiy2020image, ), CoLoR (wistuba2023, ) trains various low-rank adaptation (LoRA) (hu2021lora, ) modules for different tasks. It estimates and stores prototypes for each task and utilizes the natural clustering ability of the pre-trained model during testing to infer task IDs, selecting the corresponding LoRA component for prediction generation. In the domain of continual LLMs, architecture expansion has resurged in popularity following the rise of parameter-efficient fine-tuning (PEFT) (shazeer2017outrageously, ; aljundi2017expert, ; hu2021lora, ; dettmers2023qlora, ; lester2021power, ; li2021prefix, ), a topic we will delve into shortly (yang2024moral, ; wang2023orthogonal, ; li2024examining, ; jang2022towards, ; jin2022lifelong, ; paul2024ircoder, ; yan2023af, ; wu2024llama, ).
@@ 2.2.3. Evaluation Metrics of Continual Learning
(11) $\operatorname{OP}_{t}$ $\triangleq\tfrac{1}{t}\sum_{i=1}^{t}P_{t,i}.$
As noted in (shi2024unified, ), the OP corresponds to the primary optimization objective defined in Definition 2.4, 2.5, and 2.6. In much of the continual learning literature, once all $T$ tasks are completed, the final $\operatorname{OP}$ ($\operatorname{OP}_{T}$) is reported, with the subscript T often omitted for brevity. In some works, OP is weighted by the importance of tasks $\widetilde{\operatorname{OP}}\triangleq\tfrac{1}{T}\sum_{i=1}^{T}w_{i}P_{t,i}$, where $w_{i}=N_{i}/\sum_{j=1}^{T}N_{j}$ represents the ratio of data.
Typically, researchers report the average forgetting $F=F_{T}$ at the end of the entire training process. Forgetting quantifies the impact of learning new tasks on previously acquired knowledge. Ideally, a robust continual learning framework should achieve Backward Transfer (BWT), where learning new tasks enhances performance on prior tasks. This enhancement is typically measured by negating the forgetting, thus indicating an improvement in performance on earlier tasks.
where $b_{i}$ is the baseline performance of the model evaluated on task $i$ before undergoing continual learning. Strictly speaking, the definition of $b_{i}$ is not the same as defined in the previous work (lopez2017gradient, ; shi2024unified, ), where it is used to denote the performance of a random initialization of the model.
@@ 3. Continual Learning Meets Large Language Models: An Overview
Large language models (LLMs) are extensive in various dimensions, including the size of model parameters, pre-training datasets, computational resources, project teams, and development cycles (radford2019language, ; brown2020language, ; achiam2022chatgpt, ; achiam2023gpt, ; chowdhery2023palm, ; anil2023palm, ; touvron2023llama, ; touvron2023llama2, ). The substantial scale of LLMs presents notable challenges for development teams, particularly in keeping them updated amidst rapid environmental changes (amba2021dynamic, ; jin2022lifelong, ; dhingra2022time, ; jang2022towards, ; jang2022temporalwiki, ). To illustrate, in 2023, the average daily influx of new tweets exceeds 500 million [^1] Source: https://www.omnicoreagency.com/twitter-statistics , and training on even a subset of this large volume of data is unaffordable. Recyclable Tuning (qin2023recyclable, ) is the first work to explicitly outline the supplier-consumer structure in the modern LLM production pipeline. This structure allows us to dissect the challenges of continual LLMs from the perspectives of various roles involved. On the supplier side, the model is continually pre-trained over a sequence of large-scale unlabeled datasets. After every release of the pre-trained model, the consumer needs to utilize the stronger and more up-to-date upstream model for downstream tasks. Compared to the upstream supplier, downstream users often lack capacity of collecting and storing large-scale data, maintaining large-scale hardware systems, and training LLMs themselves. Therefore Recyclable Tuning mainly focuses on efficiently adapting an updated pre-trained LLM to downstream tasks continuously. In this survey, we further present a comprehensive framework for a modern production pipeline encompassing various studies on continual LLM pre-training, adaptation, and deployment (Fig. 1). What sets our framework in this survey apart from existing studies (wu2024continual, ) is the incorporation of two directions of continuity: Vertical Continuity and Horizontal Continuity.
@@ 3.1. Vertical Continuity (Vertical Continual Learning)
Figure 1. A high-level overview of the modern pipeline for continually pre-training and fine-tuning LLMs, where two dimensions of continuity are described. Vertical Continuity (or Vertical Continual Learning): LLM training can be vertically divided into three stages: (i) Continual Pre-Training (CPT), (ii) Domain-Adaptive Pre-training (DAP), and (iii) Continual Fine-Tuning (CFT). The main focus is the retention of the LLM’s general knowledge (prevention of vertical forgetting). Horizontal Continuity (or Horizontal Continual Learning): After the LLMs are deployed, the models are continually updated when a new set of data samples becomes available. The primary goal is to prevent horizontal forgetting in a long sequence of tasks.
Throughout the process, the unlabeled domain-specific dataset is smaller in scale than the upstream pre-training phase but larger than the final downstream task fine-tuning phase. This pattern extends to computational resources, team size, and other factors. It is important to note that vertical continuity can involve more than three stages (nijkamp2022codegen, ; lin2023geogalactica, ; rozière2024code, ; huang2023lawyer, ). In real-world applications, during domain-adaptive pre-training, additional layers can be added to accommodate multiple entities, such as various departments with distinct objectives but operating within the same domain.
Vertical Forgetting. We term the performance degradation on general knowledge of a model undergoing vertical continual learning “vertical forgetting”. As shown in Fig. 2, for vertical continual learning, the data distribution of upstream tasks partially covers the downstream, meaning the model might start off at a decent initialization for the subsequent stage of training. There are two significant challenges to be addressed to prevent vertical forgetting:
@@
Figure 2. A diagram showing two different directions of continual learning of LLMs. (a) Vertical Continual Learning of LLMs: in this case, the upstream data distribution usually partially covers the subsequent tasks’ data distribution. (b) Horizontal Continual Learning of LLMs: No constraints on the data distributions are present on horizontal continual learning. The continual LLMs need to handle the challenge of abrupt distributional shifts and longer sequence of training.
@@ 3.2. Horizontal Continuity (Horizontal Continual Learning)
Definition. Horizontal continuity (or horizontal continual learning) refers to continual adaptation across time and domains, a topic extensively explored within the continual learning community. The primary rationale for preserving horizontal continuity lies in the dynamic nature of data distribution over time. To stay updated with these content shifts, an LLM must incrementally learn newly-emerged data. Otherwise, the cost of re-training will become prohibitively expensive and impractical (chaudhry2019efficient, ; amba2021dynamic, ; su2023efficient, ; xie2023efficient, ). Empirical evidence has consistently shown that despite their impressive capabilities, LLMs struggle to generalize effectively to future unseen data, particularly in the face of temporal or domain shifts (amba2021dynamic, ; jang2022towards, ; jang2022temporalwiki, ; dhingra2022time, ). Additionally, they struggle to retain complete knowledge of past experiences when adapting to new temporal domains, although they do demonstrate a higher level of robustness against catastrophic forgetting (tao2022can, ; luo2023investigating, ; zheng2023learn, ; mehta2023empirical, ). The necessity of employing complex CL algorithms to address challenges in LLMs remains an open question. For instance, during large-scale continual pre-training, large institutions can typically afford the storage costs of retaining all historical data, rendering memory constraints not suitable. Several studies have demonstrated that with full access to historical data, simple sparse replay techniques can effectively mitigate forgetting (thengane2022clip, ; tao2022can, ; scialom2022fine, ; prabhu2023online, ; garg2024tic, ). In contrast, numerous continual learning studies have showcased superior performance compared to naive solutions, suggesting the importance of continual learning techniques in LLM training (jang2022temporalwiki, ; jin2022lifelong, ; qin2022elle, ; chen2023lifelong, ).
@@
  • Long Task Sequence. Horizontal continual learning ideally involves numerous incremental phases, particularly to accommodate temporal shifts in data distribution. A longer task sequence entails more update steps of the model, leading to inevitable forgetting of previously learned tasks. To address this challenge, researchers employ established continual learning techniques with stronger constraints, such as continual model ensemble (ramesh2021model, ).
  • Abrupt Distributional Shift. In contrast to vertical continuity, where distributional shifts are often predictable, horizontal continual learning does not impose constraints on task properties. Evidence suggests that abrupt changes in task distributions can result in significant horizontal forgetting of the model (caccia2021new, ; sarfraz2023error, ).
@@ 4. Learning Stages of Continual Large Language Models
Fig. 1 provides an overview of continually learning LLMs. Along the axis of vertical continuity, three main layers of modern continual learning emerge. The top layer, Continual Pre-Training (CPT), involves continuous pre-training of LLMs by the supplier on newly-collected data alongside existing data (Section 4.1). The middle layer, Domain-Adaptive Pre-training (DAP), prepares LLMs for domain-specific applications through additional pre-training on domain-specific unlabeled data (Section 4.2). The bottom layer, Continual Fine-Tuning (CFT), targets models for final downstream tasks on the consumer side (Section 4.3), where the model needs to be updated after deployment for the specified task.
@@ 4.1.1. CPT: Effectiveness and Efficiency
Before delving into the details of continual pre-training (CPT), it is important to address two fundamental questions: Firstly, regarding effectiveness, can CPT enhance performance on downstream tasks beyond that of the initial training on a wide range of data domains? Extensive studies have not only demonstrated the necessity of CPT for improved downstream performance (qin2022elle, ; gururangan2022demix, ; jang2022towards, ; jang2022temporalwiki, ; jin2022lifelong, ; chen2023lifelong, ), but also shown that when distributional shifts are gradual (jang2022temporalwiki, ; yildiz2024investigating, ) or somewhat correlated (gururangan2022demix, ), CPT can effectively help model generalize to unseen data. The second question is about efficiency: given the large size of LLM’s parameters and data, both old and new, can we achieve adaptation and knowledge retention in a computationally efficient way? Concerning efficiency, most studies focus on techniques for efficient knowledge retention (jin2022lifelong, ; jang2022towards, ; jang2022temporalwiki, ; li2024examining, ), which significantly overlap with the CL literature addressing catastrophic forgetting (schwarz2018progress, ; riemer2018learning, ; buzzega2020dark, ; shi2024unified, ; rebuffi2017icarl, ; ritter2018online, ; aljundi2018memory, ; rusu2016progressive, ; ramesh2021model, ; wang2022coscl, ). In contrast to prior approaches that fully utilize emergent data, some studies recognize the impracticality of this approach in real production environments. Instead, they concentrate on further improving the efficiency of adaptation. For instance, ELLE (qin2022elle, ) employs a function-preserved model expansion to facilitate efficient knowledge growth; (amba2021dynamic, ) and (xie2023efficient, ) sub-sample training data based on novelty and diversity to enhance training efficiency, achieving superior performance to full-data training. Though currently underexplored, efficient adaptation in continual pre-training is poised to become significant, given recent findings emphasizing data quality over quantity for LLM generalization (du2022glam, ; li2023quality, ; xie2024data, ; soldaini2024dolma, ).
4.1.2. General Observations on CPT
4.1.3. Distributional Shifts in CPT
This survey categorizes distributional shifts of CPT into three main types: (i) Language Shift: LLMs sequentially learn different language corpora, e.g., English $\rightarrow$ Chinese (gogoulou2024continual, ; li2024examining, ). (ii) Content Shift: LLMs sequentially learn corpora from different fields, e.g., chemistry $\rightarrow$ biology (gururangan2022demix, ; cossu2022continual, ; jin2022lifelong, ; qin2023recyclable, ; chen2023lifelong, ; gupta2023continual, ). (iii) Temporal Shift: Distributional shifts occur over time, e.g., news in 2021 $\rightarrow$ news in 2022, with a major focus on timestamp-sensitive knowledge retention and update (amba2021dynamic, ; jin2022lifelong, ; dhingra2022time, ; jang2022towards, ; jang2022temporalwiki, ). There is some work continually pre-training LLMs on datasets constructed by errors (zhao2024large, ), re-weighting the samples (chen2024take, ) or tokens (lin2024rho, ), and cannot be properly categorized, we use “Other” to represent them in Table 1.
Language Shift.(gogoulou2024continual, ) focuses on assessing LLMs’ natural ability to learn new languages sequentially (English, Norwegian, and Icelandic). With no explicit CL techniques employed for preventing horizontal forgetting, the study observes consistent positive forward transfer of the knowledge, facilitating new language acquisition regardless of the learning order. Forgetting, on the other hand, emerges as a significant challenge that cannot be mitigated by increasing LLM size. In (li2024examining, ), the degree of forgetting of previously learned language when adapting LLMs to a new language is investigated. Various CL techniques, including parameter freezing, LoRA (hu2021lora, ), and (IA)3 (liu2022few, ), are evaluated across multiple dimensions, including output language, general knowledge retention, and reliability. Preliminary experimental results highlight the non-trivial nature of addressing horizontal forgetting for CPT under the language shift. We argue that research on CPT under language shifts is in its preliminary stages for two main reasons: Firstly, the datasets’ scale, including the number of languages and total token count, remains small. Secondly, specific methods targeting language shifts have yet to be proposed; only basic combinations of existing continual learning techniques have been evaluated.
@@ 4.1.3. Distributional Shifts in CPT
Table 1. Summary of existing studies on Continual Pre-training of LLMs. The papers are organized based on their relation to CL: (i) no CL techniques are studied, (ii) CL techniques are studied as solely baselines, and (iii) new CL approaches are proposed. In the table, Dist. Shift denotes what type(s) of distributional shifts this particular study considers and is dedicated to solve. In the section of Continual Learning Tech., we mainly categorize three types of continual learning techniques that are studied in the paper: rehearsal (Rehearsal), parameter regularization (Param. Reg.), and architecture expansion (Arch. Exp.). We use “✓”, “✗”, and “♣” to denote “deployed in the proposed method”, “not studied in the paper”, and “studied as a baseline method”, respectively. It is noteworthy that we do not include naive sequential fine-tuning in this table, as it is universally studied as the important baseline method in all of the papers in the table. The papers with only “♣” (jin2022lifelong, ; jang2022temporalwiki, ; jang2022towards, ) means that no new but only existing CL techniques are studied in them, and the papers with only “✗” (gupta2023continual, ; gogoulou2024continual, ) means that no CL techniques but special aspects of fine-tuning are studied.
@@
Content Shift. Without using complex CL techniques, (yildiz2024investigating, ) explores the large-scale CPT over 159 content domains, and makes key observations about the domain structures and model properties. It shows that CPT on various domains can effectively improve models’ adaptation ability compared to DAP on single domain. Similarly, (gupta2023continual, ) continues the pre-training phase of Pythia (biderman2023pythia, ) with no complex CL techniques. The study focuses on improving CPT with simple learning rate (re)warm-up. They discover that re-warming consistently improves models trained from scratch. Built upon this simple observation, (ibrahim2024simple, ) further shows that proper combination of learning rate re-warming and re-decay, and replay of the previous data is sufficient to achieve a comparable performance as full re-training.

### v3

@@ 2.2.2. Techniques of Continual Learning
Existing CL techniques can be roughly categorized into five groups (wang2024comprehensive, ): (i) replay-based, (ii) regularization-based, (iii) architecture-based, (iv) optimization-based, and (v) representation-based. Here, we provide a concise yet comprehensive introduction to the first three categories of continual learning techniques, as they are extensively applied in continual LLMs.
@@
Replay-based methods adopt the relaxed memory constraint by keeping a small buffer of observed data and retraining the model on it when learning new tasks. Although replay-based methods may theoretically lead to loose generalization bounds (shi2024unified, ), they are valued for their simplicity, stability, and high performance, even with a small episodic memory (chaudhry2019tiny, ; riemer2018learning, ; buzzega2020dark, ; rebuffi2017icarl, ). Regularization-based methods adopt a regularization term $\lambda\left\|{\bm{\theta}}-{\bm{\theta}}_{t-1}\right\|_{\bm{\Sigma}}$ that penalizes large deviation from the history model in the parameter space, where $\|{\bm{v}}\|_{\bm{\Sigma}}={\bm{v}}^{\top}{\bm{\Sigma}}{\bm{v}}$ is the vector norm evaluated on a positive-semi-definite matrix ${\bm{\Sigma}}$, and $\lambda$ is the regularization coefficient, a hyper-parameter introduced to balance the past knowledge retention and current knowledge learning. The matrix ${\bm{\Sigma}}$ introduced is to measure the different level of importance of each parameters and their correlations in retaining the past knowledge. In practice, to reduce computational overhead, diagonal matrices are often designed to encode only the importance of each parameter (kirkpatrick2017overcoming, ; aljundi2018memory, ; rongali2021continual, ). Architecture-based methods, especially expanding the network architecture dynamically to assimilate new knowledge, is considered the most efficient form of CL (wang2022learning, ; wang2022dualprompt, ). This method primarily tackles adaptation challenges and can achieve zero-forgetting when task IDs are available during inference or can be correctly inferred (gururangan2022demix, ; wistuba2023, ). However, due to the difficulty of task ID inference, architecture expansion is predominantly utilized in TIL but is scarcely explored in DIL or CIL. In conjunction with pre-trained backbone large models like ViT (dosovitskiy2020image, ), CoLoR (wistuba2023, ) trains various low-rank adaptation (LoRA) (hu2021lora, ) modules for different tasks. It estimates and stores prototypes for each task and utilizes the natural clustering ability of the pre-trained model during testing to infer task IDs, selecting the corresponding LoRA component for prediction generation. In the domain of continual LLMs, architecture expansion has resurged in popularity following the rise of parameter-efficient fine-tuning (PEFT) (shazeer2017outrageously, ; hu2021lora, ; dettmers2023qlora, ), a topic we will delve into shortly (yang2024moral, ; wang2023orthogonal, ; li2024examining, ; jang2022towards, ; jin2022lifelong, ; paul2024ircoder, ; yan2023af, ; wu2024llama, ).
@@ 2.2.3. Evaluation Metrics of Continual Learning
There are four evaluation protocols primarily designed for continual learning. Overall Performance (OP) (ke2021achieve, ; zhang2022continual, ; zhang2023copf, ) calculates the average performance up until the current training stage, measuring the overall ability of a model balancing the performance of each task. As noted in (shi2024unified, ), OP corresponds to the primary optimization objective of continual learning, and hence receives the most attention. Forgetting (F) represents the largest performance drop observed of each task throughout the training process, averaged over all training stages. It quantifies the negative impact of learning new tasks brought to previously acquired knowledge. Ideally, a robust continual learning framework should achieve Backward Transfer (BWT), where learning new tasks enhances performance on prior tasks. BWT is measured by negating the forgetting, and hence a negative forgetting indicates a an improvement in performance on earlier tasks. Forward Transfer (FWT) measures the generalization ability of the continual learning algorithms to unseen tasks. It is defined as the difference between the current model’s performance evaluated on the future tasks and the randomly initialized model. Refer to Appendix B.1 for more details.
@@ 3. Continual Learning Meets Large Language Models: An Overview
Large language models (LLMs) are extensive in various dimensions, including the size of model parameters, pre-training datasets, computational resources, project teams, and development cycles (radford2019language, ; brown2020language, ; achiam2022chatgpt, ; achiam2023gpt, ; chowdhery2023palm, ; anil2023palm, ; touvron2023llama, ; touvron2023llama2, ). The substantial scale of LLMs presents notable challenges for development teams, particularly in keeping them updated amidst rapid environmental changes (amba2021dynamic, ; jin2022lifelong, ; dhingra2022time, ; jang2022towards, ; jang2022temporalwiki, ). To illustrate, in 2023, the average daily influx of new tweets exceeds 500 million [^1] Source: https://www.omnicoreagency.com/twitter-statistics , and training on even a subset of this large volume of data is unaffordable. Recyclable Tuning (qin2023recyclable, ) is the first work to explicitly outline the supplier-consumer structure in the modern LLM production pipeline. On the supplier side, the model is continually pre-trained over a sequence of large-scale unlabeled datasets. After every release of the pre-trained model, the consumer utilizes the stronger and more up-to-date upstream model for downstream tasks. Compared to the upstream supplier, downstream users often lack capacity of collecting and storing large-scale data, maintaining large-scale hardware systems, and training LLMs themselves. In this survey, we extend this framework and further present a comprehensive modern production pipeline encompassing various studies on continual LLM pre-training, adaptation, and deployment (Fig. 1). What sets our framework apart from existing studies (wu2024continual, ) is the incorporation of two directions of continuity: Vertical Continuity and Horizontal Continuity.
@@ 3.1. Vertical Continuity (Vertical Continual Learning)
Throughout the process, the unlabeled domain-specific dataset is smaller in scale than the upstream pre-training phase but larger than the final downstream task fine-tuning phase. This pattern extends to computational resources, team size, and other factors. It is important to note that vertical continuity can involve more than three stages (nijkamp2022codegen, ; lin2023geogalactica, ; rozière2024code, ; huang2023lawyer, ). In real-world applications, during domain-adaptive pre-training, additional “layers” can be added to accommodate multiple entities, such as various departments with distinct objectives but operating within the same domain.
Figure 1. A high-level overview of the modern pipeline for continually pre-training and fine-tuning LLMs, where two dimensions of continuity are described. Vertical Continuity (or Vertical Continual Learning): LLM training can be vertically divided into three stages: (i) Continual Pre-Training (CPT), (ii) Domain-Adaptive Pre-training (DAP), and (iii) Continual Fine-Tuning (CFT). The main focus is the retention of the LLM’s general knowledge (prevention of vertical forgetting). Horizontal Continuity (or Horizontal Continual Learning): After the LLMs are deployed, the models are continually updated when a new set of data becomes available. The primary goal is to prevent horizontal forgetting in a long sequence of tasks.
Vertical Forgetting. We term the performance degradation (in terms of general knowledge) due to vertical continual learning “vertical forgetting”. As shown in Fig. 2, for vertical continual learning, the data distribution of upstream tasks partially covers the downstream, meaning the model might start off at a decent initialization for the subsequent stage of training. Two significant challenges must be addressed to prevent vertical forgetting:
@@ 3.2. Horizontal Continuity (Horizontal Continual Learning)
Definition. Horizontal continuity (or horizontal continual learning) refers to continual adaptation across time and domains, a topic extensively explored within the continual learning community. The primary rationale for preserving horizontal continuity lies in the dynamic nature of data distribution over time. To stay updated with these content shifts, an LLM must incrementally learn newly-emerged data. Otherwise, the cost of re-training will become prohibitively expensive and impractical (chaudhry2019efficient, ; amba2021dynamic, ; su2023efficient, ; xie2023efficient, ). Empirical evidence has consistently shown that despite their impressive capabilities, LLMs struggle to generalize effectively to future unseen data, particularly in the face of temporal or domain shifts (amba2021dynamic, ; jang2022towards, ; jang2022temporalwiki, ; dhingra2022time, ). Additionally, they struggle to retain complete knowledge of past experiences when adapting to new temporal domains, although they do demonstrate a higher level of robustness against catastrophic forgetting (tao2022can, ; luo2023investigating, ; zheng2023learn, ; mehta2023empirical, ). The necessity of employing complex CL algorithms to address challenges in LLMs remains an open question. For instance, during large-scale continual pre-training, large institutions can typically afford the storage costs of retaining all historical data, rendering memory constraints meaningless. Several studies have demonstrated that with full access to historical data, simple sparse replay techniques can effectively mitigate forgetting (tao2022can, ; scialom2022fine, ; prabhu2023online, ; garg2024tic, ). In contrast, numerous continual learning studies have showcased superior performance compared to naive solutions, suggesting the importance of continual learning techniques in LLM training (jang2022temporalwiki, ; jin2022lifelong, ; qin2022elle, ; chen2023lifelong, ).
Figure 2. A diagram showing two different directions of continual learning of LLMs. (a) Vertical Continual Learning of LLMs: in this case, the upstream data distribution usually partially covers the subsequent tasks’ data distribution. (b) Horizontal Continual Learning of LLMs: No constraints on the data distributions are present on horizontal continual learning. The continual LLMs need to handle the challenge of abrupt distributional shifts and a longer sequence of training.
@@
  • Long Task Sequences. Horizontal continual learning ideally involves numerous incremental phases, particularly to accommodate temporal shifts in data distribution. A longer task sequence entails more update steps of the model, leading to inevitable forgetting of previously learned tasks. To address this challenge, researchers employ established continual learning techniques with stronger constraints, such as continual model ensemble (ramesh2021model, ).
  • Abrupt Distributional Shift. In contrast to vertical continuity, where distributional shifts are often predictable, horizontal continual learning does not impose constraints on task properties. Evidence suggests that abrupt changes in task distributions can result in significant horizontal forgetting of the model (sarfraz2023error, ).
@@ 4. Learning Stages of Continual Large Language Models
Fig. 1 provides an overview of continually learning LLMs. Along the axis of vertical continuity, three main “layers” of modern continual learning emerge. The top layer, Continual Pre-Training (CPT), involves continuous pre-training of LLMs by the supplier on newly-collected data alongside existing data (Section 4.1). The middle layer, Domain-Adaptive Pre-training (DAP), prepares LLMs for domain-specific applications through additional pre-training on domain-specific unlabeled data (Section 4.2). The bottom layer, Continual Fine-Tuning (CFT), targets models for final downstream tasks on the consumer side (Section 4.3), where the model needs to be updated after deployment for the specified task.
@@ 4.1.1. CPT: Effectiveness and Efficiency
Before delving into the details of continual pre-training (CPT), it is important to address two fundamental questions: Firstly, regarding effectiveness, can CPT enhance performance on downstream tasks beyond that of the initial training on a wide range of data domains? Extensive studies have not only demonstrated the necessity of CPT for improved downstream performance (qin2022elle, ; gururangan2022demix, ; jang2022towards, ; jang2022temporalwiki, ; jin2022lifelong, ; chen2023lifelong, ), but also shown that when distributional shifts are gradual (jang2022temporalwiki, ; yildiz2024investigating, ) or somewhat correlated (gururangan2022demix, ), CPT can effectively help model generalize to unseen data. The second question is about efficiency: given the large size of an LLM’ parameters and data, both old and new, can we achieve adaptation and knowledge retention in a computationally efficient way? Concerning efficiency, most studies focus on techniques for efficient knowledge retention (jin2022lifelong, ; jang2022towards, ; jang2022temporalwiki, ; li2024examining, ), which significantly overlap with the CL literature addressing catastrophic forgetting (schwarz2018progress, ; riemer2018learning, ; buzzega2020dark, ; shi2024unified, ; rebuffi2017icarl, ; ritter2018online, ; aljundi2018memory, ; rusu2016progressive, ; ramesh2021model, ; wang2022coscl, ). In contrast to prior approaches that fully utilize emergent data, some studies recognize the impracticality of this approach in real production environments. Instead, they concentrate on further improving the efficiency of adaptation. For instance, ELLE (qin2022elle, ) employs a function-preserved model expansion to facilitate efficient knowledge growth; (amba2021dynamic, ) and (xie2023efficient, ) sub-sample training data based on novelty and diversity to enhance training efficiency, achieving superior performance compared to full-data training. Though currently underexplored, efficient adaptation in continual pre-training is poised to become significant, given recent findings emphasizing data quality over quantity for LLM generalization (xie2024data, ; soldaini2024dolma, ).
@@
Table 1. Summary of existing studies on Continual Pre-training of LLMs. The papers are organized based on their relation to CL: (i) no CL techniques are studied, (ii) CL techniques are studied as solely baselines, and (iii) new CL approaches are proposed. In the table, Dist. Shift denotes what type(s) of distributional shifts this particular study considers and is dedicated to solve. In the section of Continual Learning Tech., we mainly categorize three types of continual learning techniques that are studied in the paper: rehearsal (Rehearsal), parameter regularization (Param. Reg.), and architecture expansion (Arch. Exp.). We use “✓”, “✗”, and “♣” to denote “deployed in the proposed method”, “not studied in the paper”, and “studied as a baseline method”, respectively. Note that we do not include naive sequential fine-tuning in this table, as it is universally studied as the important baseline method in all of the papers in the table. The papers with only “♣” (jin2022lifelong, ; jang2022temporalwiki, ; jang2022towards, ) means that only existing CL techniques are studied, without proposing new ones, and the papers with only “✗” (gupta2023continual, ; gogoulou2024continual, ) means that special aspects of fine-tuning are studied, without using CL techniques.
@@
4.1.2. General Observations on CPT
4.1.3. Distributional Shifts in CPT
This survey categorizes distributional shifts of CPT into three main types: (i) Language Shift: LLMs sequentially learn different language corpora, e.g., English $\rightarrow$ Chinese (gogoulou2024continual, ; li2024examining, ). (ii) Content Shift: LLMs sequentially learn corpora from different fields, e.g., chemistry $\rightarrow$ biology (gururangan2022demix, ; cossu2022continual, ; jin2022lifelong, ; qin2023recyclable, ; chen2023lifelong, ; gupta2023continual, ). (iii) Temporal Shift: Distributional shifts occur over time, e.g., news in 2021 $\rightarrow$ news in 2022, with a major focus on timestamp-sensitive knowledge retention and update (amba2021dynamic, ; jin2022lifelong, ; dhingra2022time, ; jang2022towards, ; jang2022temporalwiki, ).
Language Shift.(gogoulou2024continual, ) focuses on assessing LLMs’ natural ability to learn new languages sequentially. With no explicit CL techniques employed, the study observes consistent positive forward transfer of the knowledge, facilitating new language acquisition regardless of the learning order. Forgetting, on the other hand, emerges as a significant challenge that cannot be mitigated by the increasing size of LLMs. In (li2024examining, ), the degree of forgetting of previously learned language when adapting LLMs to a new language is investigated. Various CL techniques, including parameter freezing, LoRA (hu2021lora, ), and (IA)3 (liu2022few, ), are evaluated across multiple dimensions. Preliminary experimental results highlight the non-trivial nature of addressing horizontal forgetting for CPT under the language shift as well.